LTC= 150Q - 15Q^2 + 0.4Q^2. How much economic rent does the low cost producer earn?


1
Expert's answer
2021-06-22T10:19:45-0400

"LTC= 150Q - 15Q^2 + 0.4Q^2"

Long-run price=150

P=MC

"150Q - 30Q + 1.2Q^2\\30Q=1.2Q^2\\Q=25"

At this quantity the long run total cost

"150Q - 15Q^2 + 0.4Q^2=625"

Total revenue"=price\u00d7quantity"

"150×25\\=\$ 3,750"

economic rent does the low cost producer earn

=Total revenue-total cost

"=3750-625\\=\$3125"